Operation Summary for September 29, 2010 - October 1, 2010 |
Wednesday Sept. 29 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (351.6E10), with the Antiproton Source (Pbar) stacking (31.3E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with MiniBooNE off and prepared for access, with the Meson FTBF T992 and T995 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations established HEP store 8132 at 9:08 AM with an average initial luminosity (AIL) of 308.1E30. MiniBooNE came out of access at 9:44 AM. Controls personnel rebooted the Time Line Generator at 10:00 AM, which required beam operations to stop. Normal operations resumed at 10:38 AM. Operations terminated the store at 3:31 PM due to a report of a burning odor from the dump cabinet at the TeV sector B3 service building. Operators halted TeV ramp and EE Support personnel investigated the dump. During the evening shift, Operations disabled beam to experiments T992 and T995 at 6:23 PM. Operations reported at 7 PM that EE support personnel had completed their repair of the B3 dump power supply. Operations put the TeV into shot setup at 9:15 PM and established store 8135 at 10:35 PM with an AIL of 279.2E30. Operations reported that a Booster RF station had tripped off and would not reset; operators brought on the spare station. Thursday Sept. 30 The midnight shift began with Operations monitoring store 8135 (195E30) and the stash (80E10), with Pbar stacking (9.0E10), and with MiniBooNE taking beam. During the midnight shift, Operators reported that BRF6 had tripped off and needed help from Booster personnel to recover. Operations reported the loss of communication with the cryo system at TeV sector B. Operations consulted with CHL personnel and determined it could run overnight and would be further investigated in the morning. A Linac quadrupole magnet power supply tripped off and required a control card swap to fix. Operations established beam to experiments T992 and T995 at 4:09 AM. During the day shift, Operations terminated store 8135 at 1:23 PM, put the TeV into shot setup, and established store 8136 at 2:39 PM with an AIL of 300.5E30. CHL personnel rebooted two of the TeV cryo system front ends (FE) in between stores. Operations reported a breaker at the MI-50 service building had tripped off. This took down the pond water and LCW pumps at the building; it was reset and will be further investigated on Friday. During the evening shift, Operations disabled beam to experiments T992 and T995 at 5:45 PM. Operations reported that the Linac Klystron Debuncher water skid lost its temperature regulation; Linac personnel re-terminated a loose cable connector. Friday Oct. 1 The midnight shift began with Operations monitoring store 8136 (95E30) and the stash (234E10), with Pbar stacking (17E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8136 at 6:03 AM, put the TeV into shot setup, and established store 8137 at 7:09 AM with an AIL of 308.81E30. Operations reported that BRF5 tripped off and would not reset; operators bypassed the station and brought on BRF9 in its place. Operations halted the Proton Source beam at 6:36 AM in preparation for a Booster and MI access. The Booster access started at 7:35 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 10/01/2010 email Fermilab |
| Security, Privacy, Legal | |